The distribution of results from a cholesterol test has a mean of 180 and a standard deviation of 20. A sample size of 40 is drawn randomly.

Find the sum that is 1.1 standard deviations below the mean of the sums. (Round your answer to two decimal places.)
Answer #1

here expected mean of 40 drawn sample =40*180=7200

and standard deviation =20*sqrt(40)=126.4911

therefore sum that is 1.1 standard deviations below the mean of the sums =7200-1.1*126.4911=7060.86